# Learning Ordinality in Semantic Segmentation

### v1

@@ Abstract
Semantic segmentation consists of predicting a semantic label for each image pixel. Conventional deep learning models do not take advantage of ordinal relations that might exist in the domain at hand. For example, it is known that the pupil is inside the iris, and the lane markings are inside the road. Such domain knowledge can be employed as constraints to make the model more robust. The current literature on this topic has explored pixel-wise ordinal segmentation methods, which treat each pixel as an independent observation and promote ordinality in its representation. This paper proposes novel spatial ordinal segmentation methods, which take advantage of the structured image space by considering each pixel as an observation dependent on its neighborhood context to also promote ordinal spatial consistency. When evaluated with five biomedical datasets and multiple configurations of autonomous driving datasets, ordinal methods resulted in more ordinally-consistent models, with substantial improvements in ordinal metrics and some increase in the Dice coefficient. It was also shown that the incorporation of ordinal consistency results in models with better generalization abilities.
@@ 1 Introduction
Semantic segmentation, or scene parsing, is the task of attributing a semantic label to each of the pixels in an image, resulting in a segmentation map. One common problem with these deep learning segmentation models is the lack of generalization ability, which means that the network fails to make appropriate predictions when parsing a situation that did not occur in the training dataset [1]. A hypothesis is that the neural network model does not have the necessary intrinsic domain knowledge of the task – it failed to infer appropriate high-level relations from the data used to train it.
@@
Typically, ordinal problems have mostly been studied in the context of (image) classification [2, 3, 4], where the task is to classify an observation (image) as one of $\mathcal{C}_{1}\prec\mathcal{C}_{2}\prec\ldots\prec\mathcal{C}_{K}$ ordered classes (for example, the severity of a disease), as opposed to nominal classes in the case of classic nominal classification. To our knowledge, only one work [5] attempted to introduce such ordinal relations for semantic segmentation, where $\mathcal{C}_{1}\supset\mathcal{C}_{2}\supset\ldots\supset\mathcal{C}_{K}$. However, the work does not explicitly perform spatial segmentation: it would be expected that an area labelled as $\mathcal{C}_{k}$ would only have a direct boundary to the areas segmented as $\mathcal{C}_{k-1}$ and $\mathcal{C}_{k+1}$, as exemplified in Figure 1.
@@
The contributions presented in this paper can be summarized in:
  • The task of ordinal regression is formalized into ordinal consistency for each individual decision and ordinal consistency for the entire structure of the input, i.e. spatial consistency in the case of images (section 3.1).
  • Focusing on the latter case, two loss regularization terms (section 3.3) and one metric are proposed (section 3.5).
  • A thorough evaluation is performed for five biomedical datasets and an autonomous driving dataset.
@@ 2 State of the Art
Cross entropy is one of the most commonly used loss functions for image classification and segmentation problems. Defining cross entropy for a semantic segmentation problem,
@@ 2.1.1 Unimodality
The promotion of unimodality in the distribution of the model output probabilities has achieved good results in ordinal classification tasks [6, 7, 3, 8]. This can be advantageous in ordinal problems because the model should be more uncertain between ordinally adjacent classes. For example, it would not make sense for a model to output a high probability for Low and High risk of disease, but a small probability for Medium risk of disease. Figure 2 shows the difference between multimodal and unimodal distributions.
@@ 2.1.2 Ordinal Encoding
The state-of-the-art ordinal segmentation approaches have treated pixels as independent observations and promoted ordinality in their representation. However, this may be insufficient when applied to the structured image space, where pixels are dependent observations and a new level of consistency, i.e., spatial consistency, can be achieved.
@@ 3.2 Representation Consistency for Ordinal Segmentation
In ordinal segmentation, ordinal representation consistency methods encompass those methods that act on the individual pixel representation $\mathbf{x}(v)$ and decision, i.e., they impose restrictions on the pixel, taking into account its own characteristics and disregarding the context of the neighboring pixels. Such methods include the ordinal pixel encoding and consistency methods discussed in the state-of-the-art analysis [5]. The previously introduced CO2 loss (4) can be trivially adapted to segmentation by performing the regularization term for each pixel,
$\begin{split}L_{\text{O2}}(\bm{y}_{n},\hat{\bm{p}}_{n})=&\frac{1}{HW}\sum_{i=1}^{W}\sum_{j=1}^{H}\Bigg{[}\sum_{k=1}^{y_{n,i,j}}\mathrm{ReLU}(\delta+\hat{p}_{n,k,i,j}-\hat{p}_{n,k+1,i,j})\\ +&\sum_{k=y_{n,i,j}}^{K-1}\mathrm{ReLU}(\delta+\hat{p}_{n,k+1,i,j}-\hat{p}_{n,k,i,j})\Bigg{]}.\end{split}$ (8)
@@ 3.3.1 Contact Surface Loss Using Neighbor Pixels
The contact surface loss is defined by penalizing the prediction of two neighboring pixels of non-ordinally adjacent classes. This penalization term is performed by multiplying the output probabilities for neighbouring pixels,
@@ 3.4 Partially Ordered Domains
The $L_{\mathrm{O2}}^{\prime}$ term can be redefined from (8) by applying it between two neighboring classes, i.e. only when ${\ell}_{m,n}=1$, and if the edge $(m,n)$ is part of a Hasse path that includes the ground-truth label $k^{\star}$.
@@ 3.5.2 Contact Surface Metric
$\begin{split}\mathrm{CS}(\hat{\bm{y}}_{n})&=\frac{1}{2}\frac{\sum_{i=1}^{H}\sum_{j=1}^{W}\mathbbm{1}(\mathrm{CS_{dx}}(\hat{\bm{y}}_{n})_{i,j}>1)}{\sum_{i=1}^{H}\sum_{j=1}^{W}\mathbbm{1}(\mathrm{CS_{dx}}(\hat{\bm{y}}_{n})_{i,j}>0)}\\ &+\frac{1}{2}\frac{\sum_{i=1}^{H}\sum_{j=1}^{W}\mathbbm{1}(\mathrm{CS_{dy}}(\hat{\bm{y}}_{n})_{i,j}>1)}{\sum_{i=1}^{H}\sum_{j=1}^{W}\mathbbm{1}(\mathrm{CS_{dy}}(\hat{\bm{y}}_{n})_{i,j}>0)},\end{split}$ (14)
@@ 4.2 Experimental Setup
The experimental results were obtained using the UNet architecture [20] with four groups of convolution blocks (each consisting of two convolution and one pooling layers) for each of the encoder and decoder portions [^1] An open-source PyTorch implementation of the UNet architecture was used, https://github.com/milesial/Pytorch-UNet.. All datasets were normalized with a mean of 0 and a standard deviation of 1 after data augmentation, consisting of random rotation, random horizontal flips, random crops, and random brightness and contrast. The networks were optimized for a maximum of 200 epochs in the case of the biomedical datasets and a maximum of 100 epochs in the case of the autonomous driving datasets, using the Adam [21] optimizer with a learning rate of 1e-4 and a batch-size of 16. Early stopping was used with a patience of 15 epochs. After a train-test split of 80-20%, a 5-fold training strategy consisting of 4 training folds and one validation fold was applied for each training dataset. For each fold, the best-performing model on the validation dataset was selected. An NVIDIA Tensor Core A100 GPU with 40GB of RAM and an NVIDIA RTX A2000 with 12GB of RAM were used to train the networks.
The metrics used were: (1) the Dice coefficient, which evaluates the methods with respect to the ground truth labels, (2) the contact surface, which evaluates the methods with respect to the spatial ordinal consistency, and (3) the percentage of unimodal pixels metrics, which evaluates the methods with respect to the pixel-wise ordinal consistency.
@@
  • The semantic segmentation adaptation of the ordinal representation consistency $L_{\text{CO2}}$ loss function, with the imposed margin $\delta=0.05$, as recommended by the authors [3];
@@
  • The mix of representation and structural methods, through the $L_{\text{CO2}}$ + $L_{\text{CSNP}}$ loss combination of the two regularization terms.
@@ 5 Results
Tables 55 depict the main results. Since each loss term proposed is actually a penalty that is weighted with cross-entropy ($\text{loss}=\text{CE}+\lambda\,\text{term}$), each loss was evaluated using grid-search, $\lambda\in\{0.1,1,\dots,10^{4}\}$, selecting the best value for each metric.
@@
Table 5 shows that the proposed losses generally offer an improvement, as judged by the Dice coefficient. Notice that this metric is computed for each pixel, explaining why CO2 (a loss that focuses on the pixel-wise ordinal representation consistency) generally performs better. Two additional tables are presented with the proposed ordinal metrics, evaluating the consistency for the representation space (unimodality in Table 5) and the structural space (contact surface in Table 5). Naturally, CO2 again shows a very good performance for representation consistency, but the other losses start to become competitive as the structural consistency is evaluated using the contact surface metric.
Table 3: Overall results for the Dice coefficient metric.
Dataset CrossEntropy CO2 CSNP CSDT CO2+CSNP
Breast Aesthetics $93.8\pm 0.4$ $\mathbf{\hphantom{1}94.2\pm 0.5}$ $94.0\pm 0.4$ $94.0\pm 0.5$ $94.0\pm 0.2$
Cervix-MobileODT $77.0\pm 0.6$ $77.3\pm 0.7$ $\mathbf{77.4\pm 0.2}$ $76.9\pm 0.7$ $77.3\pm 0.7$
Mobbio $93.8\pm 0.1$ $\mathbf{94.1\pm 0.0}$ $93.8\pm 0.0$ $93.7\pm 0.1$ $93.8\pm 0.1$
Teeth-ISBI $74.0\pm 1.2$ $74.8\pm 0.4$ $74.9\pm 1.1$ $\mathbf{75.3\pm 1.0}$ $74.7\pm 1.0$
Teeth-UCV $90.2\pm 0.4$ $\mathbf{90.8\pm 0.3}$ $90.5\pm 0.6$ $90.3\pm 0.3$ $90.5\pm 0.3$
BDDIntersected reduced $71.2\pm 0.5$ $\mathbf{71.9\pm 0.2}$ $71.2\pm 0.6$ $70.4\pm 0.6$ $\mathbf{71.9\pm 0.2}$
BDDIntersected noabstract $37.1\pm 0.8$ $\mathbf{38.6\pm 0.8}$ $38.1\pm 0.5$ $37.2\pm 0.9$ $38.5\pm 0.9$
BDD10K $38.1\pm 1.1$ $\mathbf{39.2\pm 0.4}$ $38.2\pm 1.1$ $38.2\pm 1.2$ $\mathbf{\hphantom{1}39.2\pm 0.4}$
Cityscapes $36.0\pm 1.1$ $\mathbf{40.2\pm 0.5}$ $35.8\pm 1.1$ $36.5\pm 1.4$ $39.3\pm 0.4$
Table 4: Overall results for the percentage of unimodal pixels metric.
Dataset CrossEntropy CO2 CSNP CSDT CO2+CSNP
Breast Aesthetics $6.2\pm 0.5$ $\mathbf{99.6\pm 0.2}$ $17.0\pm 16.0$ $13.2\pm 12.3$ $56.3\pm 7.4$
Cervix-MobileODT $1.0\pm 0.2$ $\mathbf{98.9\pm 0.4}$ $1.0\pm 0.2$ $0.9\pm 0.2$ $1.2\pm 0.2$
Mobbio $0.8\pm 0.1$ $\mathbf{97.3\pm 2.3}$ $0.9\pm 0.2$ $0.7\pm 0.2$ $0.9\pm 0.0$
Teeth-ISBI $9.6\pm 3.2$ $\mathbf{97.2\pm 0.2}$ $34.7\pm 12.3$ $38.1\pm 7.1$ $93.1\pm 1.4$
Teeth-UCV $18.0\pm 1.6$ $\mathbf{98.7\pm 0.2}$ $45.7\pm 9.0$ $34.7\pm 14.7$ $81.1\pm 13.0$
BDDIntersected reduced $2.7\pm 1.7$ $\mathbf{100.0\pm 0.0}$ $39.6\pm 18.0$ $55.1\pm 12.9$ $\mathbf{100.0\pm 0.0}$
BDDIntersected noabstract $2.4\pm 1.8$ $\mathbf{100.0\pm 0.0}$ $46.0\pm 17.6$ $42.6\pm 7.1$ $\mathbf{100.0\pm 0.0}$
BDD10K $32.3\pm 5.8$ $\mathbf{100.0\pm 0.0}$ $88.8\pm 2.4$ $96.4\pm 0.5$ $\mathbf{100.0\pm 0.0}$
Cityscapes $23.0\pm 11.4$ $\mathbf{100.0\pm 0.0}$ $89.0\pm 5.2$ $97.2\pm 0.6$ $\mathbf{100.0\pm 0.0}$
Table 5: Overall results for the contact surface metric.
Dataset CrossEntropy CO2 CSNP CSDT CO2+CSNP
Breast Aesthetics $0.2\pm 0.2$ $\mathbf{0.3\pm 0.1}$ $\mathbf{0.3\pm 0.2}$ $\mathbf{0.3\pm 0.2}$ $\mathbf{0.3\pm 0.1}$
Cervix-MobileODT $\mathbf{14.5\pm 3.1}$ $\mathbf{14.5\pm 2.4}$ $14.2\pm 1.8$ $13.6\pm 1.4$ $\mathbf{14.5\pm 2.4}$
Mobbio $12.3\pm 0.4$ $\mathbf{12.5\pm 0.4}$ $12.3\pm 0.3$ $12.4\pm 0.5$ $12.0\pm 0.2$
Teeth-ISBI $30.0\pm 3.5$ $\mathbf{31.2\pm 2.6}$ $30.4\pm 1.7$ $29.1\pm 1.0$ $\mathbf{\hphantom{1}31.2\pm 2.6}$
Teeth-UCV $7.0\pm 1.1$ $6.9\pm 1.1$ $\mathbf{8.1\pm 1.3}$ $5.5\pm 0.9$ $6.9\pm 1.1$
BDDIntersected reduced $55.9\pm 3.5$ $54.9\pm 2.5$ $\mathbf{56.4\pm 1.4}$ $54.5\pm 3.2$ $52.8\pm 2.6$
BDDIntersected noabstract $47.8\pm 2.2$ $\mathbf{50.1\pm 2.2}$ $49.9\pm 1.8$ $47.0\pm 1.1$ $48.5\pm 1.5$
BDD10K $\mathbf{51.2\pm 2.1}$ $50.5\pm 2.5$ $50.3\pm 1.2$ $50.9\pm 1.4$ $48.7\pm 2.1$
Cityscapes $59.3\pm 2.0$ $\mathbf{\hphantom{1}61.4\pm 1.4}$ $58.0\pm 2.1$ $60.2\pm 2.5$ $57.6\pm 3.9$
@@
Additional results are shown in plots that depict the influence of the weight of the regularization term, $\lambda$, on each respective metric. Figures 10, 10 and 10 show, respectively, the Dice coefficient, contact surface, and unimodal metrics for the biomedical datasets. We can see that the ordinal methods successfully optimize the ordinal metrics, resulting in more ordinally consistent models. The dice coefficient metric does not improve much; however, excessive regularization can occur at some $\lambda$ values, resulting in a lower Dice metric.
@@
Figure 10 shows the Dice coefficient for the autonomous driving datasets, including the out-of-distribution domain testing. There, it can be seen that CO2 is the method with the highest impact on the generalization capability of the model, achieving a maximum absolute gain of $4.2\%$ ($11.5\%$ in relative terms) at $\lambda=10$, which means that using this loss helps the model generalize better to previously unseen scenarios. Figure 11 showcases a set of model inference outputs for this scenario, comparing cross-entropy outputs to CO2 with $\lambda=10$.

### v2

@@ Abstract
Semantic segmentation consists of predicting a semantic label for each image pixel. While existing deep learning approaches achieve high accuracy, they often overlook the ordinal relationships between classes, which can provide critical domain knowledge (e.g., the pupil lies within the iris, and lane markings are part of the road). This paper introduces novel methods for spatial ordinal segmentation that explicitly incorporate these inter-class dependencies. By treating each pixel as part of a structured image space rather than as an independent observation, we propose two regularization terms and a new metric to enforce ordinal consistency between neighboring pixels. Two loss regularization terms and one metric are proposed for structural ordinal segmentation, which penalizes predictions of non-ordinal adjacent classes. Five biomedical datasets and multiple configurations of autonomous driving datasets demonstrate the efficacy of the proposed methods. Our approach achieves improvements in ordinal metrics and enhances generalization, with up to a 15.7% relative increase in the Dice coefficient. Importantly, these benefits come without additional inference time costs. This work highlights the significance of spatial ordinal relationships in semantic segmentation and provides a foundation for further exploration in structured image representations.
@@ 1 Introduction
Semantic segmentation, or scene parsing, is the task of attributing a semantic label to each of the pixels in an image, resulting in a segmentation map. One common problem with these deep learning segmentation models is the lack of generalization ability, which means that the network fails to make appropriate predictions when parsing a situation that did not occur in the training dataset [1, 2]. While convolutional neural networks are designed to first capture low-level features from nearby pixels and progressively learn more abstract, high-level features from distant pixels [3], high-level relations are not explicitly specified. A hypothesis is that the neural network model does not have the necessary intrinsic domain knowledge of the task – it may fail to infer appropriate high-level relations from the data used to train it (e.g., the pupil being inside the iris in an eye).
@@
Typically, ordinal problems have mostly been studied in the context of (image) classification [4, 5, 6, 7], where the task is to classify an observation (image) as one of $\mathcal{C}_{1}\prec\mathcal{C}_{2}\prec\ldots\prec\mathcal{C}_{K}$ ordered classes (for example, the severity of a disease), as opposed to nominal classes in the case of classic nominal classification. DORN [8] has proposed using ordinal losses to promote order at the pixel level in-depth estimation to avoid inconsistencies where a given pixel is classified as having a high probability of being near and also a high probability of being far away but a low probability of being at a medium distance. But, to our knowledge, only one work [9] attempted to introduce ordinal relations between parts of the segmentation, where $\mathcal{C}_{1}\supset\mathcal{C}_{2}\supset\ldots\supset\mathcal{C}_{K}$. However, the work does not explicitly perform spatial segmentation: it would be expected that an area labeled as $\mathcal{C}_{k}$ would only have a direct boundary to the areas segmented as $\mathcal{C}_{k-1}$ and $\mathcal{C}_{k+1}$, as exemplified in Figure 1.
@@
While ordinal consistency has been little explored, [10] has proposed using different models for segmentation masks of classes contained inside each other. Other types of consistency from the literature include hierarchical segmentation [11], where each pixel has a taxonomy (e.g., a pixel is segmented as a small/large vehicle and then further segmented as motorcycle/bicycle if small or car/bus/truck if large). Furthermore, spatial consistency has been explored in terms of part segmentation – super-pixels have been used to impose constraints, such as the “head” appearing above “upper body” or “hair” being above “head” [12], and similarly, the function of Restricted Boltzmann Machines has been used to favor certain classes based on their spatial locations [13]. Recurrent modules have also been used to iteratively predict segmentation parts by establishing a predefined order between the parts [14]. Neither of these relates to ordinal segmentation in that sub-segmentations are contained inside super-segmentations.
The present work extends the dissertation [15] with the following contributions:
  • The task of ordinal segmentation is formalized in two consistency tasks: (i) representation consistency focuses on ordinal consistency within each individual pixel, and (ii) structural consistency focuses on ordinal consistency between structures of the input.
  • Focusing on the latter case of structural consistency, for which literature is lacking, two loss regularization terms are proposed, as well as a novel metric.
  • A thorough evaluation is performed for five biomedical datasets and an autonomous driving dataset. Methods that improve metrics of ordinal consistency are shown to also improve the Dice coefficient (a common non-ordinal semantic segmentation metric) up to 15.7% in relative terms.
@@ 2 State of the Art
Semantic segmentation, the task of assigning pixel-level labels to an image, has seen significant advancements due to the rise of convolutional neural networks. Early approaches, such as the Fully Convolutional Network (FCN) [16], replaced traditional fully connected layers with convolutional layers to enable dense pixel-wise predictions. This led to architectures like U-Net [17] and SegNet [18] that incorporated encoder-decoder structures to capture both high-level semantics and fine-grained details. Further refinements include the use of dilated (or atrous) convolutions, as introduced in DeepLab [19, 20], to capture multi-scale context without losing resolution. Another development is the incorporation of attention mechanisms, such as in PSPNet [21] and the more recent Transformer-based approaches, e.g. ones based on Swin [22], which have improved performance by better capturing long-range dependencies.
2.1 Loss Functions
In these works, a common loss is cross entropy, albeit auxiliary losses may also be used, such as a differentiable version of the Dice coefficient. Defining cross entropy for a semantic segmentation problem,
@@ 2.2.1 Unimodality
The promotion of unimodality in the distribution of the model output probabilities has achieved good results in ordinal classification tasks [23, 24, 5, 25]. This can be advantageous in ordinal problems because the model should be more uncertain between ordinally adjacent classes. For example, it would not make sense for a model to output a high probability for Low and High risk of disease but a small probability for Medium risk of disease. Figure 2 shows the difference between multimodal and unimodal distributions.
@@ 2.2.2 Ordinal Encoding
The state-of-the-art ordinal segmentation approaches have treated pixels as independent observations and promoted ordinality in their representation. However, this may be insufficient when applied to the structured image space, where pixels are dependent observations.
@@ 3.2 Representation Consistency for Ordinal Segmentation
In ordinal segmentation, ordinal representation consistency methods encompass those methods that act on the individual pixel representation $\mathbf{x}(v)$ and decision, i.e., they impose restrictions on the pixel, taking into account its own characteristics and disregarding the context of the neighboring pixels. Such methods include the ordinal pixel encoding and consistency methods discussed in the state-of-the-art analysis [9]. The previously introduced O2 term (4) can be trivially adapted to segmentation by performing the regularization term for each pixel,
$\begin{split}L_{\text{O2}}(\bm{y}_{n},\hat{\bm{p}}_{n})=&\frac{1}{HW}\sum_{i=1}^{W}\sum_{j=1}^{H}\Bigg{[}\\ &\sum_{k,k^{\prime}\in\mathcal{S}^{A}(y_{n,i,j})}\mathrm{ReLU}(\delta+\hat{p}_{n,k,i,j}-\hat{p}_{n,k^{\prime},i,j})+\\ &\sum_{k,k^{\prime}\in\mathcal{S}^{D}(y_{n,i,j})}\mathrm{ReLU}(\delta+\hat{p}_{n,k,i,j}-\hat{p}_{n,k^{\prime},i,j})\Bigg{]}.\end{split}$ (8)
where $k,k^{\prime}\in\mathcal{S}^{A}(y_{n,i,j})$ and $k,k^{\prime}\in\mathcal{S}^{D}(y_{n,i,j})$ are pairs of classes that have an ascending and descending order, respectively, which typically would be $S^{A}(y_{n,i,j})=\{(k,k+1)\mid k<y_{n,i,j}\}$ and $S^{D}(y_{n,i,j})=\{(k+1,k)\mid k>y_{n,i,j}\}$.
@@ 3.3.1 Contact Surface Loss Using Neighbor Pixels
The contact surface loss is defined by penalizing the prediction of two neighboring pixels of non-ordinally adjacent classes. This penalization term is performed by multiplying the output probabilities for neighboring pixels,
@@ 3.4 Partially Ordered Domains
The $L_{\mathrm{O2}}^{\prime}$ term can be redefined from (8) by applying set $\mathcal{S}$ between two neighboring classes whenever ${\ell}_{m,n}=1$, and if the edge $(m,n)$ is part of a Hasse path that includes the ground-truth label $k^{\star}$.
@@ 3.5.2 Contact Surface Metric
$\begin{split}\mathrm{CS}(\hat{\bm{y}}_{n})&=\frac{1}{2}\frac{\sum_{i=1}^{H}\sum_{j=1}^{W}\mathbbm{1}(\mathrm{CS_{dx}}(\hat{\bm{y}}_{n})_{i,j}\geq 2)}{\sum_{i=1}^{H}\sum_{j=1}^{W}\mathbbm{1}(\mathrm{CS_{dx}}(\hat{\bm{y}}_{n})_{i,j}\geq 1)}\\ &+\frac{1}{2}\frac{\sum_{i=1}^{H}\sum_{j=1}^{W}\mathbbm{1}(\mathrm{CS_{dy}}(\hat{\bm{y}}_{n})_{i,j}\geq 2)}{\sum_{i=1}^{H}\sum_{j=1}^{W}\mathbbm{1}(\mathrm{CS_{dy}}(\hat{\bm{y}}_{n})_{i,j}\geq 1)},\end{split}$ (14)
@@ 4.2 Experimental Setup
Following the same procedure as [9], experimental results were obtained using the UNet architecture [17] with four groups of convolution blocks (each consisting of two convolution and one pooling layers) for each of the encoder and decoder portions [^1] An open-source PyTorch implementation of the UNet architecture was used, https://github.com/milesial/Pytorch-UNet.. All datasets were normalized with a mean of 0 and a standard deviation of 1 after data augmentation, consisting of random rotation, random horizontal flips, random crops, and random brightness and contrast. The networks were optimized for a maximum of 200 epochs in the case of the biomedical datasets and a maximum of 100 epochs in the case of the autonomous driving datasets, using the Adam optimizer with a learning rate of 1e-4 and a batch-size of 16. Early stopping was used with a patience of 15 epochs. After a train-test split of 80-20%, a 5-fold training strategy consisting of 4 training folds and one validation fold was applied for each training dataset. For each fold, the best-performing model on the validation dataset was selected. An NVIDIA Tensor Core A100 GPU with 40GB of RAM and an NVIDIA RTX A2000 with 12GB of RAM were used to train the networks.
The metrics used were: (1) the Dice coefficient, which evaluates the methods with respect to the ground truth labels, (2) the contact surface, which evaluates the methods with respect to the structural ordinal consistency, and (3) the percentage of unimodal pixels metrics, which evaluates the methods with respect to the pixel-wise ordinal consistency.
@@
  • The semantic segmentation adaptation of the ordinal representation consistency $L_{\text{O2}}$ loss function, with the imposed margin $\delta=0.05$, as recommended by the authors [5];
@@
  • The mix of representation and structural methods, through $L_{\text{CE}}$ + $L_{\text{O2}}$ + $L_{\text{CSNP}}$ loss combination of the two regularization terms.
@@ 5 Results
Tables 55 depict the main results. Since each loss term proposed is actually a penalty that is weighted with cross-entropy ($\text{loss}=\text{CE}+\lambda\,\text{term}$), each loss was evaluated using grid-search, $\lambda\in\{0.1,1,\dots,10^{4}\}$, selecting the best $\lambda$ according to the Dice metric. In the tables, bold indicates that the method has produced significantly better results than the baseline (just CE) using a one-tailed paired t-test with alpha=0.05.
@@
Table 5 shows that the proposed losses generally offer an improvement, as judged by the Dice coefficient. Notice that this metric is computed for each pixel, explaining why CE+O2 (a regularization term that focuses on the pixel-wise ordinal representation consistency) generally performs better. Two additional tables are presented with the proposed ordinal metrics, evaluating the consistency for the representation space (unimodality in Table 5) and the structural space (contact surface in Table 5). Naturally, CE+O2 again shows a very good performance for representation consistency, but the other losses start to become competitive as the structural consistency is evaluated using the contact surface metric.
Table 3: Overall results for the Dice coefficient metric (%), a common metric for semantic segmentation. Introducing ordinal knowledge into the loss function improves this metric. Higher is better.
Dataset CE CE+O2 CE+CSNP CE+CSDT CE+O2+CSNP
Breast Aesthetics $93.8\pm 0.5$ $94.2\pm 0.6$ $94.0\pm 0.5$ $94.0\pm 0.4$ $94.0\pm 0.4$
Cervix-MobileODT $77.0\pm 0.7$ $77.2\pm 0.4$ $77.4\pm 0.2$ $76.9\pm 0.7$ $77.3\pm 0.7$
Mobbio $93.8\pm 0.1$ $\mathbf{94.1\pm 0.0}$ $93.8\pm 0.0$ $93.7\pm 0.2$ $93.8\pm 0.1$
Teeth-ISBI $74.0\pm 1.6$ $74.8\pm 0.4$ $74.9\pm 1.4$ $\mathbf{75.3\pm 1.3}$ $74.7\pm 1.1$
Teeth-UCV $90.2\pm 0.5$ $90.8\pm 0.4$ $90.5\pm 0.3$ $90.3\pm 0.3$ $\mathbf{90.5\pm 0.4}$
BDDIntersected reduced $71.2\pm 0.5$ $71.0\pm 0.4$ $71.2\pm 0.6$ $70.4\pm 0.8$ $71.9\pm 0.2$
BDDIntersected noabstract $37.1\pm 1.0$ $\mathbf{38.6\pm 1.0}$ $38.1\pm 0.7$ $37.2\pm 1.1$ $\mathbf{38.5\pm 1.1}$
BDD10K $38.1\pm 1.1$ $39.1\pm 0.5$ $38.2\pm 1.2$ $38.2\pm 1.2$ $39.2\pm 0.6$
Cityscapes $36.0\pm 1.3$ $\mathbf{40.2\pm 0.7}$ $35.8\pm 1.2$ $36.5\pm 1.4$ $\mathbf{39.3\pm 0.4}$
Table 4: Overall results for the percentage of unimodal pixels metric (%). This is a representation metric that focuses on the ordinal consistency probabilities produced for each individual pixel. CE+O2, which also focuses on individual pixels, is better at addressing this metric. Higher is better.
Dataset CE CE+O2 CE+CSNP CE+CSDT CE+O2+CSNP
Breast Aesthetics $6.2\pm 0.7$ $8.1\pm 2.6$ $5.9\pm 2.1$ $6.2\pm 1.5$ $6.7\pm 0.9$
Cervix-MobileODT $1.0\pm 0.2$ $\mathbf{98.9\pm 0.4}$ $0.5\pm 0.2$ $0.9\pm 0.3$ $1.2\pm 0.3$
Mobbio $0.8\pm 0.1$ $\mathbf{97.3\pm 2.6}$ $0.9\pm 0.2$ $0.7\pm 0.2$ $\mathbf{0.9\pm 0.0}$
Teeth-ISBI $9.6\pm 3.4$ $\mathbf{67.1\pm 3.8}$ $\mathbf{11.2\pm 4.6}$ $\mathbf{13.5\pm 3.4}$ $\mathbf{15.5\pm 4.4}$
Teeth-UCV $18.0\pm 1.8$ $\mathbf{96.8\pm 1.0}$ $25.7\pm 6.8$ $19.7\pm 4.1$ $\mathbf{73.3\pm 6.6}$
BDDIntersected reduced $2.7\pm 1.8$ $3.4\pm 2.0$ $\mathbf{18.8\pm 9.4}$ $1.6\pm 1.3$ $6.4\pm 4.0$
BDDIntersected noabstract $2.4\pm 1.9$ $\mathbf{11.7\pm 4.1}$ $3.7\pm 2.8$ $6.8\pm 5.2$ $\mathbf{12.7\pm 6.1}$
BDD10K $32.3\pm 6.4$ $36.1\pm 9.2$ $31.7\pm 1.7$ $34.4\pm 2.7$ $\mathbf{49.9\pm 9.3}$
Cityscapes $23.0\pm 12.0$ $\mathbf{69.9\pm 5.4}$ $38.1\pm 3.9$ $24.9\pm 2.4$ $\mathbf{88.4\pm 4.0}$
Table 5: Overall results for the contact surface metric (%). This is a structural metric that measures consistency between adjacent pixels. In this case, the proposed terms (CSNP and CSDT) improve the ordinal differentiation between neighbors. Lower is better.
Dataset CE CE+O2 CE+CSNP CE+CSDT CE+O2+CSNP
Breast Aesthetics $0.2\pm 0.2$ $0.2\pm 0.2$ $0.1\pm 0.1$ $0.2\pm 0.2$ $0.3\pm 0.2$
Cervix-MobileODT $14.5\pm 3.3$ $\mathbf{1.1\pm 0.4}$ $10.8\pm 2.7$ $13.6\pm 1.7$ $14.5\pm 2.5$
Mobbio $12.3\pm 0.5$ $\mathbf{4.2\pm 0.2}$ $12.3\pm 0.4$ $12.4\pm 0.6$ $\mathbf{11.5\pm 0.3}$
Teeth-ISBI $30.0\pm 3.9$ $26.1\pm 2.4$ $28.4\pm 1.4$ $27.1\pm 1.8$ $27.5\pm 1.2$
Teeth-UCV $7.0\pm 1.4$ $\mathbf{3.0\pm 0.7}$ $\mathbf{2.3\pm 1.0}$ $5.5\pm 1.0$ $\mathbf{1.8\pm 0.5}$
BDDIntersected reduced $55.9\pm 4.0$ $54.9\pm 2.9$ $\mathbf{49.1\pm 2.3}$ $54.5\pm 3.8$ $52.8\pm 3.0$
BDDIntersected noabstract $47.8\pm 2.6$ $47.8\pm 2.1$ $47.8\pm 2.6$ $46.6\pm 1.6$ $47.9\pm 2.0$
BDD10K $51.2\pm 2.2$ $50.5\pm 2.8$ $50.3\pm 1.4$ $50.9\pm 1.5$ $\mathbf{48.4\pm 2.3}$
Cityscapes $59.3\pm 2.2$ $61.4\pm 1.8$ $56.9\pm 3.7$ $60.2\pm 2.8$ $\mathbf{55.5\pm 2.2}$
(As bold when the method is better than the baseline (CE) using a one-tailed paired t-test with alpha=0.05.)
Figure 6: Comparison of the effect of changing the $\lambda$ coefficient controlling the strength of the CSDT loss term (Mobbio dataset) with (a) Ground-truth, (b) $\lambda=0.1$, (c) $\lambda=1$, (d) $\lambda=10$. The latter (d) is an example of excessive regularization, where the model includes a sclera border (dark gray) between the background (black) and the iris (light gray), which does not exist in the ground truth.
@@
Additional results are shown in plots that depict the influence of the weight of the regularization term, $\lambda$, on each respective metric. Figures 8, 9 and 10 show, respectively, the Dice coefficient, contact surface, and unimodal metrics for the biomedical datasets. We can see that the ordinal methods successfully optimize the ordinal metrics, resulting in more ordinally consistent models. The dice coefficient metric does not improve much; however, excessive regularization can occur at some $\lambda$ values, resulting in a lower Dice metric. An example of this over-regularization is presented in Figure 6.
@@
Figure 11 shows the Dice coefficient for the autonomous driving datasets, including the out-of-distribution domain testing. There, it can be seen that CE+O2 is the method with the highest impact on the generalization capability of the model, achieving a maximum absolute gain of $4.2\%$ ($11.5\%$ in relative terms) at $\lambda=10$, which means that using this loss helps the model generalize better to previously unseen scenarios. Figure 12 showcases a set of model inference outputs for this scenario, comparing cross-entropy outputs to CE+$\lambda$ O2 with $\lambda=10$.